# for Dynamic Neural Radiance Fields

### v3

@@ IV EVALUATION
ReFusion [13] StaticFusion [41] DynaSLAM (G) [23] DynaSLAM (N+G) [23] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.068
balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.038
balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.036
balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.028
crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.061
crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.056
crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.055
kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020
kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017
moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014
moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.026
moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.167
moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176
person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.024
person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.035
placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.018
placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.020
placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022
placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.117
removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015
removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.021
removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.181
synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.006
synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.011
Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.051/0.181
IV-A Datasets
To assess DynaMoN, we compare both the camera localization and the novel view synthesis part with state-of-the-art approaches on two challenging datasets.
TUM RGB-D - Dynamic subset [12]
The TUM RGB-D dataset has a resolution of $640\times 480$. While the dataset mainly focuses on static parts, five sequences were rated slightly dynamic and four highly dynamic.
BONN RGB-D Dynamic [13]
The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences. The number of recordings per scene varies. The images have a resolution of $640\times 480$.
For the novel view synthesis results, we follow the usual splitting using every eighth frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D Dynamic dataset (bottom). Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
@@ IV-A Datasets
TABLE III: Novel view synthesis results on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that the camera pose could not be regressed.
@@
Sequences RoDynRF
HexPlane
@@
fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.83 / 0.725
fr3/sitting-static 9.67 / 0.123 - 28.55 / 0.899 28.95 / 0.903
fr3/sitting-xyz 10.10 / 0.150 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
fr3/sitting-rpy 8.35 / 0.121 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
fr3/sitting-halfsphere 7.46 / 0.157 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
fr3/walking-static 10.01 / 0.159 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
fr3/walking-xyz 10.02 / 0.171 - 24.05 / 0.797 24.34 / 0.796
fr3/walking-rpy 8.81 / 0.201 - 24.18 / 0.790 24.70 / 0.796
fr3/walking-halfsphere 8.79 / 0.205 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
Mean 8.84 / 0.157 - 25.05 / 0.809 25.60 / 0.819
TABLE IV: Novel view synthesis results on BONN RGB-D Dynamic [13]. The - denotes that the camera pose could not be regressed.
@@
IV-B Implementation Details
For semantic masking, we use DeepLabV3 [42] with a ResNet50-backbone [43] trained on the large-scale COCO dataset [44] and use human masks. For motion masking, we use the trained model of DytanVO [35]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence. For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold of 60 % for the maximum amount of pixels that contain motion. Consequently, masks exceeding this threshold are discarded to ensure that there are always enough pixels for the dense bundle adjustment. Additionally, we only consider a pixel as dynamic if the confidence of the network exceeds 0.95 for the first and 0.98 for the remaining mask refinement steps.
While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1 024. We employ the default optimization settings for the dynamic NeRF from [4] and use a multi-step learning rate decay scheduler for the pose refinement. The initial learning rate for the scheduler is set to $1*10^{-5}$ at the beginning of each pose optimization cycle and decayed every 1 000 iterations by a factor of $0.9$. As described beforehand, the actual used learning rate of the optimizer is retrieved by multiplication with the percentage of static pixels in the image. The number of optimization iterations is set to 200 000, unless otherwise stated.
To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth. Besides the COLMAP comparison, we compare our approach with RoDynRF [34]. For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM. We follow their training guideline and do not employ prior pose initialization.
TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamic crowd2 scene with sparse input trajectory.
@@ IV-B Implementation Details
TABLE VI: Analysis of the runtime on the TUM RGB-D fr2/desk-person scene. $\ddagger$ means 100 000 training iterations and a batch size of 512.
$\dagger$ means 100 000 training iterations and a batch size of 1 024. The time format is hours:minutes:seconds.
@@
DynaMoN (Ours) 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
DynaMoN (Ours) $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
DynaMoN (Ours) $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
IV-C Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS). Looking at the mean value of the RMSE of the translational trajectory error over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios. Moreover, comparing the RMSE of the rotational ATE, see Figure 3, our MS&SS-model achieves lower values than DROID-SLAM. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for both rotation and translation.
IV-D Novel View Synthesis Quality
To evaluate our approach from a NeRF perspective we consider two aspects. First, the traditional way of generating camera ground truth when using NeRF is the use of COLMAP. Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP. Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses. Third, we compare our approach with RoDynRF [34] which regresses the camera pose along with a dynamic NeRF.
The results for the TUM RGB-D dataset are denoted in Table III. DynaMoN leads to a higher novel view synthesis quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses. Qualitative examples of our renderings are shown in Fig. 4.
On the BONN RGB-D Dynamic dataset, COLMAP is challenged on several sequences (cf. Tab. IV). DynaMoN can retrieve camera poses for all scenes (Tab. II) and performs well on novel view synthesis for all sequences (Tab. IV). In addition, the PSNR is increased by the alternating training scheme. However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results. Examples of renderings can be seen in Fig. 4.
IV-E Robustness on Sparse Trajectories
In this section, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory. Using only every tenth frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality. Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
IV-F Runtime and Memory Evaluation
Finally, the runtime of DynaMoN is measured and the memory requirements are mentioned. Both are presented in comparison with RoDynRF [34] on a workstation with an NVIDIA RTX 4090 graphics card with 24 gigabytes of video random-access memory (VRAM). We report the results for the TUM RGB-D fr2/desk-person sequence in Table VI as it is by far the longest scene of both datasets comprising 4 067 frames. In terms of memory requirement, RoDynRF can only be trained with a batch size of 512 in this setting, while DynaMoN is more memory-efficient, enabling a batch size of 1 024. From the experiments, it can be concluded that the total time for pre-processing, training and rendering of DynaMoN is comparable to RoDynRF if the default training strategy with 200 000 iterations is utilized. However, if our model is trained for 100 000 iterations which is the default value of RoDynRF, the total time can be greatly reduced without a significant drop in photometric accuracy. For this shorter training, the NeRF and pose optimization of DynaMoN alternate every 2 500 iterations. Furthermore, a decrease of the batch size to the same value as in the competitor method further speeds up the training time at the cost of a longer rendering time while staying comparably accurate. It has to be noted that the pre-processing of RoDynRF consisting of monocular depth prediction, optical flow prediction, motion mask generation and dataloading takes about eight times longer than our pre-processing loading the data and generating initial poses as well as motion masks.
@@ V DISCUSSION
In comparison to traditional sparse reconstructions [15, 14, 16], our approach addresses camera localization and dynamic novel view synthesis in combination. While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11]. Nevertheless, compared to combined dynamic camera localization and dynamic novel view synthesis approaches [34], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
Moreover, our approach shows an improved performance compared to the state-of-the-art for camera localization. Only using our motion masks already improves the performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results.
The results presented in Table III indicate that RoDynRF [34], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [34] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.
VI CONCLUSION
We present DynaMoN, a motion-aware fast and robust camera localization approach for novel view synthesis. DynaMoN can handle not only the motion of known objects using semantic segmentation masks but also that of unknown objects using a motion segmentation mask. Furthermore, it retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation. In addition, DynaMoN is trained with a novel iterative training scheme that refines NeRF parameters and the initially provided poses in an alternating fashion. Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for novel view synthesis on scenes with high camera motion.
ACKNOWLEDGMENT
We thank d.hip for providing a campus stipend.

### v4

@@ IV-A Datasets
DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [43] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (SS) OURS (MS)
[44] [20] [22] [43] [45] [38] [23] [19]
fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.028 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.036 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.017/0.036 0.016/0.039
TABLE III: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [15]. RGB-D results are reported by Palazzolo et al. [15]. For DROID-SLAM and ours, we calculated the results.
Sequences RGB-D RGB
ReFusion [15] StaticFusion [46] DynaSLAM (G) [47] DynaSLAM (N+G) [47] DROID-SLAM [19] OURS (MS&SS) OURS (SS) OURS (MS)
balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.030 0.068
balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.027 0.038
balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.037 0.036
balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.029 0.028
crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.033 0.061
crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.029 0.056
crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.052 0.055
kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020 0.020
kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017 0.017
moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014 0.014
moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.027 0.026
moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.152 0.167
moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176 0.176
person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.033 0.024
person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.023 0.035
placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.042 0.018
placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.019 0.020
placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022 0.022
placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.173 0.117
removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015 0.015
removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.020 0.021
removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.179 0.181
synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.007 0.006
synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.007 0.011
Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.049/0.179 0.051/0.181
IV-B Implementation Details
For semantic masking, we use DeepLabV3 [48]/ResNet50 [49] trained on COCO [50] to obtain human masks. For motion masking, we apply DytanVO [38]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence.
When the camera motion exceeds that of the DytanVO training sequences, the motion segmentation module produces more false positives. To counter this, we set a threshold of 60% for the maximum number of motion pixels; masks exceeding this threshold are discarded to maintain sufficient pixels for dense bundle adjustment. Additionally, a pixel is considered dynamic only if the confidence is above 0.95 for the first mask refinement step and 0.98 for subsequent steps.
For NVS, we use the full image size. During camera localization, the images are downsampled by half to remain consistent with existing trajectory experiments [19].
We set the loss parameter $\lambda_{TV}$ to 0.005 and sample $\lvert\mathcal{R}\rvert=1,024$ rays per batch. We use the default optimization settings for the dynamic NeRF [4], with a multi-step learning rate decay scheduler for pose refinement. The initial learning rate is $1\times 10^{-5}$, decayed by a factor of 0.9 every $1,000$ iterations and reset in each pose optimization cycle. The learning rate is scaled by the percentage of static pixels in the image. The number of optimization iterations is set to $200,000$, unless specified otherwise.
IV-C Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table II and Table III. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS), using only semantic masks (SS) as well as using both types of masks in combination (MS&SS). Looking at the mean value of the RMSE of the translational ATE over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for translation. Moreover, comparing the RRA, see Table I, our MS&SS-model achieves higher values than DROID-SLAM on the BONN RGB-D dataset, while all methods perform equally well on TUM RGB-D. It can be noted that using one type of mask is usually sufficient to improve over DROID-SLAM. In some cases, like the person-tracking scene, where the combination performs worse than SS or MS, the SLAM backbone seems to have too less pixels left for the bundle adjustment.
IV-D Novel View Synthesis Quality
To evaluate the NVS quality of our approach, we consider several aspects. First, we compare it with the traditional way of generating camera ground truth for 4D NeRF [4] and 4D GS [32] using COLMAP, see Tables IV and V. We follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth. Second, we compare our approach with RoDynRF [37] without initial COLMAP poses, shown in Table IV. We follow RoDynRF’s training guidelines. Third, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses, see Tables  IV, Table V and  VI. Qualitative examples of renderings on both datasets can be found in Figure 3.
On both datasets, COLMAP is challenged by the dynamics, and training with COLMAP camera poses can only be performed on a subset of scenes. DynaMoN retrieves all camera poses for both datasets, see Tables II and III. As visualized in Figure 3, the visual results of COLMAP-based NeRF are noisy on both datasets. DynaMoN’s camera poses can show improved results. The complete DynaMoN can further improve details. In terms of PSNR/SSIM, DynaMoN leads to a higher NVS quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses for either 4D NeRF or 4D GS, see Table IV.
@@ IV-D Novel View Synthesis Quality
TABLE IV: NVS results on TUM RGB-D. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that the camera pose could not be regressed. The * denotes random point cloud initialization, as the one from COLMAP lead to no convergence. We report PSNR/SSIM.
@@
Sequences RoDynRF [37]
4D Gaussians [32]
HexPlane [4]
@@
fr2/desk-person 6.05 / 0.127 - - 23.04 / 0.680 24.83 / 0.725
fr3/sitting-static 9.67 / 0.123 - - 28.55 / 0.899 28.95 / 0.903
fr3/sitting-xyz 10.10 / 0.150 26.40 / 0.886 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
fr3/sitting-rpy 8.35 / 0.121 24.15 / 0.814 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
fr3/sitting-halfsphere 7.46 / 0.157 24.42 / 0.830 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
fr3/walking-static 10.01 / 0.159 16.39 / 0.575* 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
fr3/walking-xyz 10.02 / 0.171 - - 24.05 / 0.797 24.34 / 0.796
fr3/walking-rpy 8.81 / 0.201 - - 24.18 / 0.790 24.70 / 0.796
fr3/walking-halfsphere 8.79 / 0.205 21.72 / 0.746 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
Mean 8.84 / 0.157 - - 25.05 / 0.809 25.60 / 0.819
TABLE V: NVS results on BONN RGB-D Dynamic. The - denotes that the camera pose could not be regressed. We report PSNR/SSIM.
@@
TABLE VI: Analysis of different pose retrieval steps on NVS quality, evaluated on the crowd2 scene of the BONN RGB-D Dynamic dataset. We assess robustness using a sparse input trajectory (every 10th frame).
@@
IV-E Robustness on Sparse Trajectories
Dynamic scenes pose more challenges regarding the robustness of an approach compared to static environments. In Table VI, we showcase the robustness of our approach. We use only every tenth frame of the crowd2 sequence from the BONN RGB-D Dynamic dataset. As reported in Table VI, we show how the underlying camera localization approach influences the NVS result. When comparing DROID-SLAM camera poses to the ones from our DynaMoN backbone and our neural camera pose refinement, the best NVS result can be achieved using DynaMoN. Looking at the PSNR values in Table VI, it can be noted that each of our proposed components increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
IV-F Runtime and Memory Evaluation
We compare the runtime and memory consumption of DynaMoN with RoDynRF [37] on a workstation with an NVIDIA RTX 4090 graphics card with 24 GB of VRAM. We report the results for the TUM RGB-D fr2/desk-person sequence in Table VII as it is by far the longest scene of both datasets comprising 4 067 frames. Using the COLMAP configuration from Instant-NGP, COLMAP runs for 38:28:16 (hours:minutes:seconds), see Table VII. In terms of memory requirement, RoDynRF can only be trained with a batch size of 512 in this setting, while DynaMoN is more memory-efficient, enabling a batch size of 1 024. From the experiments, it can be concluded that the total time for pre-processing, training and rendering of DynaMoN is comparable to RoDynRF if the default training strategy with 200 000 iterations is utilized. However, if our model is trained for 100 000 iterations which is the default value of RoDynRF, the total time can be greatly reduced without a significant drop in photometric accuracy. For this shorter training, the NeRF and pose optimization of DynaMoN alternate every 2 500 iterations. Furthermore, a decrease of the batch size to the same value as in the competitor method, RoDynRF, further speeds up the training time at the cost of a longer rendering time while staying comparably accurate.
It has to be noted that the pre-processing of RoDynRF consisting of monocular depth prediction, optical flow prediction, motion mask generation and dataloading takes about eight times longer than our pre-processing loading the data and generating initial poses as well as motion masks.
Figure 3: Qualitative Results for NVS on TUM RGB-D (top) and BONN RGB-D Dynamic (bottom). Ground Truth (left), novel views views from HexPlane + COLMAP (2nd column), improved camera poses from our backbone (3rd column), and our full model (right).
TABLE VII: Analysis of the runtime on TUM RGB-D fr2/desk-person. Ours refers to 200,000 iterations / 1,024 batch size. The symbol $\ddagger$ indicates 100,000 iterations / 512 batch size, while $\dagger$ represents 100,000 iterations / 1.024 batch size. Time is given in the format hours:minutes:seconds
@@ IV-F Runtime and Memory Evaluation
Ours 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
Ours $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
Ours $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
@@ V DISCUSSION
In comparison to traditional sparse reconstructions [18, 17, 19], our approach addresses camera localization and dynamic NVS in combination. While this leads to more pleasing 3D visualizations and NVS results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11]. Nevertheless, compared to combined dynamic camera localization and dynamic NVS approaches [37], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
Moreover, our approach shows an improved performance compared to the state-of-the-art for camera localization. Only using our motion masks already improves the performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results. The results presented in Table IV indicate that RoDynRF [37], despite being a state-of-the-art method that integrates combined camera regression with NVS, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [37] and HexPlane [4] employing COLMAP poses on the BONN RGB-D Dynamic and TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table VI, and achieves better performance in NVS on both datasets.
Looking at the results of 4D GS with COLMAP poses, combining robust camera localization and 4D GS seems to be a promising future research direction.
VI LIMITATIONS
DynaMoN is constrained by the scene representation capabilities of dynamic NeRFs. Although it performs well on large-scale datasets with up to over $4,000$ images, handling scenes with close to $8,000$ frames introduces significant challenges to the camera pose localization [51] and the NeRF [52]. Nevertheless, our method works effectively with datasets like TUM RGB-D and BONN RGB-D Dynamic. Still, larger datasets [53] pose challenges, for NeRF and the camera localization [51]. The underlying dynamic NeRF in this work has difficulties representing environments of this scale, limiting its applicability to more extensive datasets.
VII CONCLUSION
We present DynaMoN, a motion-aware, fast and robust camera localization approach for NVS. DynaMoN can cope with the motion of objects of known and unknown classes using the combined segmentation and motion mask resulting in the dynamics mask. DynaMoN retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation. In addition, DynaMoN is trained with a novel iterative training scheme that refines NeRF parameters and the initially provided poses in an alternating fashion. Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for NVS on scenes with high camera motion.